how do solve x square root3= square root 12
Found 2 solutions by ilana, fractalier:
Answer by ilana(307)

well sqrt12=sqrt4*sqrt3, so x=sqrt4=2.
We also could have said x=sqrt12/sqrt3, so x=sqrt(12/3), so x=sqrt4=2.
So x=2.

Answer by fractalier(6550)

Divide both sides by root 3...
x root 3 = root 12
(x root 3)/(root 3) = (root 12)(root 3)
x = root 4
x = 2